[ Use similar triangles to answer the following questions.
A tree casts a shadow 6 ft long. At the same time a vertical rod 4 ft high casts a shadow that is 3 ft long. How tall is the tree? ]

Answer by TimothyLamb(4379)

h/6 = 4/3
h = 6(4/3)
h = 2(4)
---
h = 8 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php